OpenStudy (valeriav096):

When solving negative one over five(x − 25) = 7, what is the correct sequence of operations?

OpenStudy (seratul):

So \[\frac{ 1 }{ 5(x-25) } = 7\] ?

OpenStudy (danjs):

\[\large \frac{ -1 }{ 5*(x-25) } = 7\] or \[\frac{ -1 }{ 5 }*(x-25)=7\]

OpenStudy (valeriav096):

\[-\frac{ 1 }{ 5}(x − 25) = 7\]

OpenStudy (valeriav096):

like this.

OpenStudy (danjs):

ok.. to get rid of the fraction, you can multiply both sides of the equation by 5 first

OpenStudy (danjs):

-1*(x - 25) = 35

OpenStudy (danjs):

distribute the -1 into the parenthesis -1*x + (-1)*(-25) = 35 -x + 25 = 35 right?

OpenStudy (valeriav096):

then we subtract 25 from both sides?

OpenStudy (danjs):

yes you can -x = 10 divide by -1